OpenStudy (anonymous):

4. A certain freely falling object, released from rest, requires 1.50 s to travel the last 30.0 m before it hit the ground. a. Find the velocity of the object when it is 30.0 m above the ground. b. Find the total distance the object travels during the fall.

OpenStudy (anonymous):

tell me about the equations that comes in your mind ?

OpenStudy (anonymous):

\[S = u t + \frac{ 1 }{ 2 } a t ^{2}\]

OpenStudy (anonymous):

try to complete it now

OpenStudy (anonymous):

\[v^{2}-v_{o}^{2}=2gh\]

OpenStudy (anonymous):

ok i will try

OpenStudy (anonymous):

i can't do it !!!

OpenStudy (anonymous):

see equation vo=0,then.............

OpenStudy (anonymous):

s= 30m u = ? ( initial velocity of the object ) a = 9.81 m/s^2 ( accn of free fall ) t = 1.5 s s = ut + 1/2 at^2 \[u = \frac{ S - 1/2 a t^2 }{ t }\] \[u = \frac{ 30 - ( 0.5 \times 9.81 \times 1.5^2) }{ 1.5 } \] \[u = 12.6 m/s\]

OpenStudy (anonymous):

for part ii) v = u + at v = at because u = 0 since the object is being released and v is the speed calculated above t = v/a

OpenStudy (anonymous):

s = ut + 1/2 a t^2 u = 0 , s = 1/2 a t^2 \[s = \frac{ 1 }{ 2 } \times a \times t ^{2}\] \[s = \frac{ 1 }{ 2 } \times 9.81 \times \left( \frac{ 12.6 }{ 9.81 } \right)^{2}\] \[s = 8.0917...\] \[therfore total distance = 8.0917 + 30 = 38.0917.. = 38.1 m \]